OpenStudy (anonymous):

Solve the equation by factoring: x2 - 3x - 28 = 0 can you show steps please?

OpenStudy (anonymous):

x2squared

OpenStudy (anonymous):

x2 -3x -28= 0 x2 -3x = 28 (divide by 2) x-3x=14 -2x=14 (divide by 2) x=-7

OpenStudy (anonymous):

Find two factors of -28 that when multiplied get -28, BUT when added also equal -3. So it would be, x^2 -7x + 4x -28. Then factor then, it will look like this, (x^2 - 7x) (4x-28) = 0. Factor out the common factor... x(x-7) +4(x-7) =0. Next, you "combine" like terms and you put x and +4 together to get (x+4)(x-7) =0. Then you "make two new equations" to look like this, x+4=0 and x-7=0. Finally, solve for x... x=-4 and x=7